Find the two repeated limits of the function f(x,y)=(y-x/y+x)(1+x^2/1+y^2) at (0,0).Does the simultaneous limit of f exist as (x,y) to (0,0)?Give reasons for your answer
1
Expert's answer
2020-03-16T13:40:23-0400

The given function is ,

"f(x,y)= \frac{(y-x)}{(y+x)}×\frac{(1+x^2)}{(1+y^2)}"

Now, we have to find the repeated limit of "f(x,y)"

"\lim_x\to0" ("\lim_y\to0 f(x,y)" )"=\lim_x\to0" "(" -1×("1+x^2)" ")"

"=-1" .


"\lim_y\to0" "(" "\lim_x\to0f(x,y)" ")" "=\lim_y\to0" "(" "\frac{1}{(1+y^2)}" ")"

"=1" .

No,simultaneous limit of f at (0,0) does not exist.

Because the above repeated limit are not equal.